INAF Science

01:17
  • OSCO
  • Missed the star on the first attempt.
  • Didn't restart the w_unit processes after Guido re-started the adsec.
  • w_stop_all/start_all fixed it right up.
1:46
  • On to ESM target.
  • g_tt_0.15_ho1_0.30_ho2_0.30.fits
  • gives slightly better slope rms than LNgains.
3:10
  • Heading to next ESM target.
  • It's the spectroscopic one from last night.
  • Had to do a pointing check.
  • The seeing has degraded...
  • 2.24"......
04:45
  • Still chugging along.

-- GregoryTaylor - 15 Oct 2020
Topic revision: r4 - 16 Oct 2020, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback